OpenStudy (anonymous):

a. Suppose a market research company finds that at a price of p = $20, they would sell x = 42 tiles each month. If they lower the price to p = $10, then more people would purchase the tile, and they can expect to sell x = 52 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint: Write an equation using two points in the form (x,p).

OpenStudy (saifoo.khan):

just a sec.

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

which appendix is that?

OpenStudy (anonymous):

appendix c

OpenStudy (saifoo.khan):

The two points are: (42,20) and (52,10) Slope: (10-20)/(52-42) Slope= -10/10 Slope= -1 Substituting the values: y = mx+b 20 = -1(42) + b b = 62 so: p(x) = -x + 62

OpenStudy (anonymous):

is that the answer i put or do it have to be in (x,p) form?

OpenStudy (saifoo.khan):

It says: a. Write your answer in the form p = mx + b

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (saifoo.khan):

No problem! ^_^

OpenStudy (anonymous):

so i'm saying p(x)=-x+62 is my final answer after all of that?

OpenStudy (saifoo.khan):

are u like those people who dont get algebra at all?

OpenStudy (saifoo.khan):

Oh Yes.

OpenStudy (anonymous):

i am....most of it I dont get for real...but i'm trying though

OpenStudy (saifoo.khan):

lol, alright.. best of luck with your coming weeks.

OpenStudy (anonymous):

I have a feeling you'll be hearing from me again...LOL

OpenStudy (anonymous):

Thanks again

OpenStudy (saifoo.khan):

Haha, Sure. the best part, i got the same feeling as well. LOL.

OpenStudy (anonymous):

Good one!!

OpenStudy (saifoo.khan):

;]

OpenStudy (saifoo.khan):

how's the situation?